THE INFRARED SPECTRUM OF CH_{5}^{+}

\underline{TAKESHI OKA}, Department of Chemistry, Department of Astronomy and Astrophysics, and the Enrico Fermi Institute, The University of Chicago, Chicago, IL 60637.

Protonated methane, CH_{5}^{+}, has unusual C-H bondings and is a new prototype of spectroscopic specimen. While all of the five protons are strongly bound to the central carbon atom with well defined C-H stretch potentials, angles between C-H bonds are highly fluxional. Ab initio theory predicts lowest energy for an ”equilibrium structure” with C_{s} symmetry but the barriers separating the 24 equivalent structures are extremely low and almost non-existent when zero point vibrations are taken into account.^{a}^{b} We have identified nearly 1000 spectral lines of CH_{5}^{+} in a liquid N_{2}-cooled hydrogen dominated plasma using a gas mixture of H_{2}:CH_{4} ∼ 100:1.^{c} The spectral lines are weak and do not show obvious symmetry or regular pattern. The identification required extensive studies of the spectra of other carbocations CH_{3}^{+} (and CH_{2}^{+}) and C_{2}H_{3}^{+} (and C_{2}H_{2}^{+}) whose spectra appear in the same wavelength region with much higher intensities. Spectroscopy and plasma chemistry of those carbocations will be discussed and the raw observed spectrum of CH_{5}^{+} will be presented without assignment nor even qualitative understanding.

aP. R. Schreiner et al., J.Chem.Phys. \underline{99}, 3716 (1993)

bH. Mueller et al., J.Chem.Phys. \underline{106}, 1863 (1997)

cE. T. White, J. Tang and T. Oka, Science \underline{284}, 135 (1999)